Opened 11 years ago

Last modified 7 years ago

#169 assigned flight processing

GB08/08, flight day 175/2008, Wytham Woods

Reported by: mggr Owned by: adbe
Priority: alpha 4 high Milestone: Reprocessing
Component: Processing: general Keywords:
Cc: jatt Other processors:

Description

Data location: ~arsf/arsf_data/in_progress/2008/flight_data/uk/GB08_08-2008_175_Wytham_Woods/

Data arrived from ARSF via network transfer (~18hrs) finishing 24/June/2008.

Scientific details: see ~/arsf/arsf_data/in_progress/2008/ARSF_Applications/GB_and_NL/GB08-08 Science Case.pdf

PI: R Harding

Sensors:

  • ATM
  • Eagle
  • Hawk

Note only applanix processing required, then forward to Andrew Wilson/CEH.

Quicklook

TBD.

Change History (8)

comment:1 Changed 11 years ago by mggr

Note only applanix processing required, then forward to Andrew Wilson/CEH.

comment:2 Changed 11 years ago by benj

  • Cc jatt added

Applannix processed by jatt, rsynced to ~arsf/arsf_data/in_progress/2008/flight_data/uk/GB08_08-2008_175_Wytham_Woods

comment:3 Changed 11 years ago by mark1

Data has been delivered on the 7th July 2008

comment:4 Changed 10 years ago by mggr

  • Resolution set to fixed
  • Status changed from new to closed

Data sent to NEODC for archiving.

comment:5 Changed 7 years ago by benj

  • Resolution fixed deleted
  • Status changed from closed to reopened

Level 1 data doesn't seem to have gotten from AKW to NEODC. Eagle/Hawk need to be reprocessed and sent to NEODC, please notify Charles George (ctg@…) when done.

comment:6 Changed 7 years ago by benj

  • Owner set to adbe
  • Status changed from reopened to assigned

comment:7 Changed 7 years ago by adbe

Created Delivery

comment:8 Changed 7 years ago by adbe

Ready to deliver at:

/users/rsg/arsf/workspace/GB08_08-2008_175_Wytham_Woods/processing/delivery/GB08_08-175-hyperspectral-20120201

Note: See TracTickets for help on using tickets.